Heidi has 12 more $5 bills than $2 bills. altogether she has $137 how many of each kind does she have?
Answer by mananth(16946)

$2 bills be x
$ 5 bills = x+12
..
Total value = $137
..
2x+5(x+12)=137
2x+5x+60=137
7x+60=137
subtract 60 from both sides
7x=137-60
7x=77
/7
x= 11 , the number of $2 bills
$ 5 bills will be x+12 = 11+12 =23
...
m.ananth@hotmail.ca